(b) What intensity level of the horns sound is dB him and 476 Hz when the train is receding from him. Using these frequendies, he calculates the speed of the train. What value does he find? (Assume the speed of sound in air is 343 89
Answer #1

4)
First find the intensity of a 56dB sound using β = 10dB*log(I/10-12)

so I = 10(56/10)*10-12= 3.981*10-7 W/m2

Now P = I*A = 3.981*10-7*4π*(10000m)2 = 500.26 W

b) at 46m the intensity = P/A = 500.26W/(4π*462) = 0.01881 W/m2

This corresponds to a level of 10dB*log(0.01881/10-12) = 102.74dB

5)
With the train approaching at speed, the observed frequency is 345 m 345 m/s 481Hz- 345 s 345 m/s -345 m/s-v, As the train recedes, the observed frequency is sa 345 m/s+0 345 m/s-(-D)345 m/s+v. 416 Hz Dividing equation (1) by (2) gives

487 345+vt 476 345 - vt


(487 - 476) / (487 +476) ={ 345 + v - (345 - v) } / {345 + v + 345 - v}

11/ 963 = 2v / 2(345)
11/ 963 = v / 345
vt = 3.94 m/s